# HAMMER: Heterogeneous, Multi-Robot Semantic Gaussian Splatting

### v1

@@ IV-B Multi-robot 3DGS Mapping
A naive approach to multi-robot map building is to collect all robot data into a large dataset, align the data through bulk SfM, and optimize a $3$DGS model from the static dataset. This procedure is very similar to how neural fields are conventionally trained. However, offline optimization of the map prevents the robot team from using the map during deployment, reducing the map’s effectiveness.
@@
Instead, after each robot is aligned to the global reference frame, any data from that robot that was received by the mapping server is fused into the collaborative 3DGS map. To achieve this functionality, our $3$DGS mapping system performs online optimization on a stream of data and flexibly incorporates data from heterogeneous devices with possibly different camera sensors and geometric scene data (e.g. point clouds or depth images). When only RGB images are available, monocular depth estimation [29] could be used as geometric scene data if it produces metric depth. However, handling such cases is outside the scope of our implementation of HAMMER. In the following, we outline HAMMER’s multi-robot $3$DGS mapping module.
IV-B1 Representation
@@
3DGS models the opacity and color of the environment using explicit Gaussian primitives, which are optimized based on a differentiable, tile-based rasterization process into a photometric loss. The optimizable parameters of each Gaussian include a mean $\mu\in\mathbb{R}^{3}$, a covariance $\Sigma\in S^{3}_{++}$ (the geometric extent of the Gaussian), an opacity $\alpha\in[0,1]$, and an RGB color $c\in[0,1]^{3}$. The color is further represented as a set of learned spherical harmonics coefficients, yielding view-dependent color to model realistic visual appearance.
@@
Furthermore, we seek to model semantics in the environment. Specifically, HAMMER encodes the per-pixel semantic embedding, extracted from 2D color images using CLIP [30, 31], into the $3$DGS. Some works [32] augment the parameters of each Gaussian with a semantic channel and regress it directly from a image reconstruction loss. However, these semantic embeddings can range from 500 to 1000 dimensions, which is prohibitively expensive to inference and store. Therefore, an additional encoding network is used to compress the CLIP embedding. This encoding network is learned in conjunction with the 3DGS model, but can be difficult to optimize while streaming data.
@@
Instead, HAMMER optimizes a continuous feature field, parameterized by a neural network with hash-grid positional encodings. This feature field $\phi_{s}(x):\mathbb{R}^{3}\rightarrow\mathbb{R}^{d_{s}}$ maps positions in the $3$DGS frame $\mathcal{T}^{g}$ to a feature vector of dimension $d_{s}$. To query the semantic embedding of a Gaussian, we evaluate $\phi_{s}$ at the mean. During training, we re-project the $3$DGS rendered depth and ground-truth CLIP images into $3$D, producing a labeled point cloud. The point cloud is then directly used to supervise the feature field.
@@
A key challenge when mapping with data streams from heterogeneous robots is compensation for each sensor’s differing image signal processing (ISP) pipelines (e.g. exposure and white-balancing). These pipelines can have a significant impact on the appearance of the resulting color images. Consequently, these appearance shifts can produce view inconsistency in the color between devices, despite their content being the same. The view inconsistency can cause erroneous, non-existent structures in the 3DGS map, called floaters, due to the optimizer reconciling contradicting signals from images captured from devices with different ISP pipelines. To avoid this undesirable behavior, we utilize the bilateral grid-based ISP compensation method proposed in [33]. Specifically, for each image, we create additional bilateral grid parameters that compensate for the appearance differences between devices by matching the color of the rendered images to the ground-truth images through affine transformations.
IV-B2 Online Optimization
@@
IV-B3 Pose Refinement
During online map optimization, HAMMER refines not only the individual camera poses, but also the relative transformation between each robot’s local frame and the global frame. Specifically, we optimize an $SE(3)$ offset from the aligned camera pose to the refined camera pose. Regularization on both the rotation and translation is used to avoid catastrophic drift in the poses. We use a similar approach for fine-tuning the alignment of each of the robot’s local frame $\mathcal{T}^{j}$ to the global frame $\mathcal{T}^{g}$.
V Experiments
We validate HAMMER on two real world scenes with data from a heterogeneous team of robots, and show that it outperforms baselines and is effective for semantic navigation.
Figure 4: Incremental map quality for HAMMER and baselines in two scenes. HAMMER (Basement: 18.07, Home: 18.9) outperforms Di-NeRF* (15.8 and 16.3) in map quality, demonstrating that accurate robot alignment method is essential for mapping. It also outperforms the Individuals baselines (i.e. no collaboration), highlighting the benefits of collaborative map building. Oracle serves as an upper bound on map quality.
V-A Experimental Setup
V-A1 Implementation
HAMMER uses ROS2 [34] to stream data from all robots to a server. Software for HAMMER is based off of a modified version of NerfBridge [35]. The server is a desktop equipped with an AMD Ryzen 7 5800x CPU, a NVIDIA RTX 4090 GPU, and 32 GBs of RAM.
The heterogeneous teams are composed of Turtle Bots (non-holonomic ground robots equipped with stereo cameras) and humans wearing Aria glasses [1] that are equipped with a broad range of onboard sensors. The ground robots are outfitted with a forward facing ZED Mini stereo camera roughly 0.5m above the ground. The ZED camera is connected to a NVIDIA Jetson Orin Nano that handles on-board sensor data processing. The camera uses its proprietary SLAM system integrated with the ZED SDK for pose estimation. The ground robot is connected to the server using 5GHz WiFi, and streams the estimated pose, 1280x720 MJPEG compressed color images, and 1280x720 depth images at 10Hz.
Aria glasses provide their own onboard compute and pose estimate using the Aria Project’s proprietary SLAM system. The glasses produce 1440x1440 fisheye color images and semi-dense point clouds at 10Hz. Unlike the ZED, the Aria devices do not have a ROS interface, so we convert data recorded from these devices into ROS bags and play these bags back during experiments to simulate streaming. For all robots, we assume access to intrinsic camera calibration information.
V-A2 Trial Details
We test HAMMER in two real-world environments, which we label Home and Basement. Home is a multi-room apartment and contains regions of high feature richness (open rooms with plants/furniture) and low richness (hallways and rooms with blank walls). The scene also includes regions of drastic lighting changes due to open windows with outside natural light, resulting in wide variations in color image appearances due to camera exposure compensation. In Home, we deploy two Aria glasses and one ground robot.
Basement is composed of a large lab space, kitchenette, long hallways, and a conference room. Lighting conditions in this environment are more uniform due to the presence of continuous overhead lighting, but also includes several regions with sparse visual features like hallways and reflective floor-to-ceiling glass windows. Two Aria glasses and two ground robots are deployed in Basement.
Evaluation data is produced by holding out some of the data streamed from each device during the trials. To assess the accuracy of HAMMER’s pose alignment and refinement, we process all data with COLMAP (i.e. bulk SfM) to get accurate pose estimates. This dataset serves as an upper bound on the quality of the aligned streamed dataset. The bulk SfM procedure takes about 2-5 hours for each trial, illustrating the need for HAMMER’s real-time alignment module.
V-A3 Baselines
We compare HAMMER against three baselines. The first is Oracle, which uses the output of the bulk SfM procedure to optimize a $3$DGS map. Again, Oracle provides an upper bound on the performance of HAMMER because the SfM poses are treated as ground truth in our evaluation set. Next, Individuals optimizes independent $3$DGS maps using each robot’s own data, which forms a set of baselines. Importantly, to ensure a fair comparison, Individuals uses the same data collected by executing HAMMER, but it has an extra advantage in that the 3DGS map is trained in bulk with all data available at once (from a single robot). Individuals shows the best quality possible from a single robot, thereby highlighting HAMMER’s added value in using multi-device collaboration to improve map quality.
To our knowledge, the only comparable work to HAMMER is Di-NeRF [12], which proposes collaboratively optimizing a NeRF while using the same gradient-based alignment correction scheme discussed in Section IV-B3. However, rather than solving for initial relative frame transformations like HAMMER, Di-NeRF initializes translation and rotation to zero and identity, respectively. HAMMER has additional differences to Di-NeRF, which we modify to provide a fair comparison. First, Di-NeRF is a distributed optimization algorithm, while HAMMER is centralized. Therefore, we solve the centralized problem proposed by Di-NeRF which forms an upper bound on the performance of Di-NeRF. Second, Di-NeRF does not perform online training, so we use cached data from HAMMER trials for map optimization. Again, this modification gives Di-NeRF an advantage compared to HAMMER because it has access to all of the data at the start of training. Finally, Di-NeRF uses a NeRF representation, which is typically less performant than $3$DGS. Therefore, we replace the NeRF with a $3$DGS, retaining the hyper-parameters of HAMMER when possible. We refer to this baseline as Di-NeRF*.
@@ V-A3 Baselines
V-B Results
In both trials, HAMMER is able to successfully construct a quality semantic $3$DGS map. To evaluate HAMMER, we compare its map reconstruction quality to the baselines and include run time analysis to illustrate that HAMMER can be used during robot deployments.
V-B1 Map Reconstruction
@@
Figure 3 visualizes meshes extracted from the $3$DGS maps constructed by HAMMER as they evolve with time during the online map optimization process. Embedded in the meshes are trajectories taken by each robot during deployment, highlighting where each robot contributed data and the overall spatial distribution of data. Qualitatively, HAMMER demonstrates superior photo-realistic renders on the evaluation set (Figure 5) compared to Di-NeRF*, which fails to resolve robot alignments and therefore cannot accurately match the ground-truth images. HAMMER also approaches the quality of the Oracle baseline.
@@
We quantify reconstruction quality based on the average peak signal to noise ratio (PSNR) of each method on the held-out evaluation data set (Figure 4). HAMMER dramatically outperforms Di-NeRF*, suggesting that the gradient-based alignment of Di-NeRF fails to converge to accurate inter-robot alignments. The set of Individuals baselines underperforms HAMMER as expected, highlighting the benefit of robot collaboration in map building. Note that the performance of some individual maps is relatively high due to good performance on their respective evaluation sets, but degrade drastically when benchmarked on another robot’s evaluation set. Finally, as expected, Oracle is an upper bound on the performance of HAMMER but is significantly slower.
V-B2 Runtime
@@
HAMMER runtime is dominated by the $3$DGS map optimizer and the inter-robot alignment module. Processing a single data-frame in the map optimizer ranges between 15 ms (66 Hz) when the map is sparsely populated to 38 ms (26 Hz). Slower iterations typically occur immediately after a robot has been aligned and all of its cached data is consolidated into the map. The runtime of the alignment module is dominated (90% - 99.5% of the runtime) by the localized SfM procedure. Each SfM alignment requires about 33.6s to solve. The remaining runtime is primarily the feature-matching based verification for proposed image correspondences (approximately 2.6s).
V-B3 Multi-robot Semantic Motion Planning
@@
We demonstrate the utility of HAMMER for collaborative downstream tasks by completing a multi-robot motion planning task with language-specified goals in Basement using a HAMMER $3$DGS map. To set a goal location, we query an object’s location using natural language. Then, using CLIP, we extract the semantic embedding for the goal query and search the HAMMER $3$DGS map for the most semantically similar Gaussian using cosine similarity. Next, we sample the surrounding space for a nearby location in free space, which we set as the goal location. Once a goal is specified, we use Splat-Nav [7] to plan a collision-free trajectory from the initial position of each robot to the goal location. Figure 6 shows optimized trajectories for each of the four robots in Basement with 9 semantically-specified goals. Searching the map for each semantic goal location required about 20ms, and the subsequent trajectories were computed within 0.2s to 0.75s depending on the distance to the goal. All robots successfully completed the task.
Figure 6: Motion plans of robots navigating to language-specified goals with trajectories from Splat-Nav [7].
VI Conclusion and Limitations

### v2

@@ III-C Multi-robot 3DGS Mapping
A naive approach to multi-robot map building is to collect all robot data into a large dataset, align the data through bulk SfM, and optimize a $3$DGS model from the static dataset. This procedure is similar to how neural fields are conventionally trained. However, offline optimization of the map prevents the robot team from using the map during deployment and on-demand, reducing the map’s effectiveness.
@@
Instead, HAMMER continually trains the global map from steaming robot data, adding new robots and devices to the set of contributing robots with the alignment module from Sec. III-B. To achieve this functionality, our $3$DGS mapping system performs online optimization on a stream of data and flexibly incorporates data from heterogeneous devices with possibly different camera sensors and geometric scene data (e.g. point clouds or depth images). When only RGB images are available, monocular depth estimation [30] could be used as geometric scene data if it produces metric depth. However, handling such cases is outside the scope of our implementation of HAMMER.
Figure 3: Meshes extracted from HAMMER’s map during runtime (left) and the final map reconstruction (right). The attribution meshes show the part of the map contributed by each robot: Aria 1 (red), Aria 2 (green), GR 1 (orange), and GR 2 (blue).
III-C1 Representation
@@
3DGS models the opacity and color of the environment using explicit Gaussian primitives, which are optimized based on a differentiable, tile-based rasterization process into a photometric loss. The optimizable parameters of the $k$-th Gaussian include a mean $\mu_{k}\in\mathbb{R}^{3}$, a covariance $\Sigma_{k}\in S^{3}_{++}$ (the geometric extent of the Gaussian), an opacity $\alpha_{k}\in[0,1]$, and an RGB color $c_{k}\in[0,1]^{3}$. The color is further represented as a set of learned spherical harmonics coefficients, yielding view-dependent color to model realistic visual appearance.
@@
Aside from color, HAMMER is also able to model semantics in the environment. Specifically, per-pixel semantic embeddings, extracted from color images using CLIP [31, 32], are distilled into the $3$DGS. The choice of pixel-wise semantic feature extractors is flexible, but we choose CLIP because of its strong zero-shot performance [33]. Some works [34] augment the parameters of each Gaussian with a semantic channel and regress it directly from an image reconstruction loss. However, these semantic embeddings can vary up to 1000 dimensions, which is prohibitively expensive to inference and store. Therefore, some methods store an encoding network used to compress the CLIP embedding. This encoding network is learned in conjunction with the $3$DGS, but can be difficult to optimize while streaming data.
@@
Alternatively, HAMMER optimizes a continuous feature field, parameterized by a neural network with hash-grid positional encodings. This feature field $\phi_{s}(x):\mathbb{R}^{3}\rightarrow\mathbb{R}^{d_{s}}$ maps positions in the $3$DGS frame $\mathcal{T}^{g}$ to a feature vector of dimension $d_{s}$. To query the semantic embedding of a Gaussian, we evaluate $\phi_{s}$ at $\mu_{k}$. During training, we re-project the $3$DGS rendered depth and ground-truth CLIP images into $3$D, producing a labeled point cloud. The point cloud is then directly used to supervise the feature field. Gradient flow from the point cloud to the underlying model is blocked in order to simplify gradient computation as we do not observe improved scene reconstruction otherwise.
@@
A key challenge when mapping with heterogeneous data streams is compensation for each sensor’s different image signal processing (ISP) pipelines (e.g. exposure and white-balancing). These pipelines can have a significant impact on the appearance of the resulting color images. Consequently, these appearance shifts can produce view inconsistency in the color between devices, despite their content being the same. The view inconsistency can cause erroneous, non-existent structures in the 3DGS map, called floaters, due to the optimizer reconciling contradicting signals from images captured from devices with different ISP pipelines. To avoid this undesirable behavior, we utilize the bilateral grid-based ISP compensation method proposed in [35]. Specifically, for each image, we create additional bilateral grid parameters that compensate for the appearance differences between devices by matching the color of the rendered images to the ground-truth images through affine transformations.
III-C2 Online Optimization
@@
III-C3 Pose Refinement
Although the alignment module produces robust estimates of the local-to-world transforms, it cannot account for gradual drift or other temporal noise. To combat this deficiency, in addition to optimizing the scene parameters, HAMMER treats the SLAM pose and the local-to-world transform $\mathcal{T}^{j}\to\mathcal{T}^{g}$ as initial estimates that are continually optimized as part of $3$DGS training, incrementally deforming each agent’s trajectory to create a visually consistent global map. Specifically, we optimize an $SE(3)$ offset from the aligned camera pose to the refined camera pose. Regularization on both the rotation and translation is used to avoid catastrophic drift in the poses. We use a similar approach for fine-tuning $\mathcal{T}^{j}\to\mathcal{T}^{g}$.
IV Experiments
First, we compare HAMMER to state-of-the-art baselines [14, 13] by assessing their reconstruction accuracy on the ReplicaMultiAgent dataset [15, 14]. However, ReplicaMultiAgent only contains scenes from simulated environments, and lacks heterogeneous robots/sensing devices and challenging real-world scene conditions (e.g. motion blur, diverse lighting). Therefore, to showcase the generalizability of HAMMER and its real-time deployment in real-world environments, we also assess its performance in two different hardware trials with data collected using real robots.
IV-A Experimental Setup
IV-A1 Implementation
Unlike other baselines, HAMMER is fully integrated into the ROS2 ecosystem [36] to stream data from all robots to a server, and uses a modified version of NerfBridge [37] for map building. The mapping server is a desktop computer with a NVIDIA RTX 4090 GPU. 5GHz WiFi is used to communicate during hardware trials.
ReplicaMultiAgent is composed of two robots covering four scenes featuring mesh reconstructions of real indoor environments. Each robot produces RGBD images. To simulate real-time deployment within the environments, we create 2 minute long ROSBags that stream data at 20Hz. Each robot uses KISS-ICP [38] to perform localization and streams this data simultaneously with the other robot.
[…]
@@ IV-A1 Implementation
IV-A2 Scene Details
The MultiAgentReplica dataset contains four indoor scenes: Off-0, Apt-0, Apt-1, and Apt-2. Scenes have one to three rooms with no lighting variation.
Our hardware trials consist of two real-world scenes: Home and Basement. Home is a 6 room apartment and contains regions of high feature richness (open rooms with plants/furniture) and low richness (hallways and rooms with blank walls). The scene also includes regions of drastic lighting changes due to open windows with outside natural light, resulting in wide variations in color image appearances due to camera exposure compensation. We deploy two Aria glasses and one ground robot in this scene. Basement is composed of a large lab space, kitchenette, long hallways, and a conference room. This environment includes several regions with sparse visual features like hallways and reflective floor-to-ceiling glass windows. Here, two Aria glasses and two ground robots are deployed.
IV-A3 Baselines
In MultiAgentReplica, we benchmark HAMMER against CP-SLAM [14] and MAGiC-SLAM [13]. However, importantly, neither of the public implementations of the baselines are real-time capable. In the real-world trials, we compare HAMMER to three different baselines. The first is Oracle, which uses bulk SfM on the same sensor data used by HAMMER to optimize a $3$DGS map. Oracle provides an upper bound on the performance of HAMMER because the bulk SfM poses serve as a “ground truth” in our evaluation set. Next, Individuals is a set of baselines, where each corresponds to a single robot optimizing its own independent $3$DGS maps from its own data. Because each robot cannot map the entire scene, any single map is not generalizable, highlighting the need for a global map.
Because both MAGiC-SLAM and CP-SLAM are not real-time capable, do not have publicly available code for integration with hardware/ROS, and do not support heterogeneous devices, we cannot compare these methods to HAMMER in the real-world trials. Instead we compare to Di-NeRF [16], which proposes collaboratively optimizing a NeRF using gradient-based robot frame alignment. We modify Di-NeRF to provide a fair comparison to HAMMER by centralizing its map optimization (previously distributed), training it on the data cached during HAMMER’s runtime, and replacing its NeRF with a $3$DGS. We refer to this baseline as Di-NeRF*.
IV-B Results
IV-B1 MultiAgentReplica
In Table I, we compare the scene reconstructions performance of HAMMER to MAGiC-SLAM and CP-SLAM. HAMMER outperforms both baselines on all averaged metrics, and does so at least 25$\times$ faster than CP-SLAM and 16$\times$ faster than MAGiC-SLAM. The key advantage of HAMMER over these baselines is that it decouples local robot pose estimation from the joint mapping process, leading to significantly faster runtimes. Superior reconstruction results also demonstrate that this decoupling does not degrade scene reconstruction.
IV-B2 Hardware Trials
@@
HAMMER is able to successfully construct quality $3$DGS maps, visualized in (Fig. 3) as the 3DGS evolves in time during the online map optimization process. Embedded in the figure are the trajectories traversed by the robots, along with colors indicating where each robot contributed data. Qualitatively, HAMMER demonstrates superior novel-view renders on the evaluation set (Fig. 4) compared to Di-NeRF*, which fails to resolve robot alignments and therefore cannot accurately match the ground-truth images. HAMMER also approaches the quality of Oracle.
@@
We quantify reconstruction quality based on the average PSNR of each method on the held-out evaluation dataset (Fig. 5), comprised of 10 held-out frames per device. HAMMER dramatically outperforms Di-NeRF* which fails to converge to accurate inter-robot alignments. Individuals underperforms HAMMER since evaluation views span the entire scene, highlighting the benefit of robot collaboration in map building. Note that the performance of some individual maps is relatively high due to good coverage of the entire scene from particular robots. Finally, Oracle upper bounds the performance of HAMMER due to access to “ground-truth” poses. It also has access to the entire dataset at optimization time, leading to more optimal reconstructions compared to those generated from streaming data like in HAMMER. However, Oracle is much slower to compute.
Figure 5: Map quality over time for HAMMER and baselines in two scenes. HAMMER outperforms Di-NeRF*, demonstrating the necessity of accurate robot alignment. It also outperforms Individuals, highlighting the benefits of collaboration. Oracle serves as an upper bound on map quality.
IV-B3 Runtime
@@
HAMMER runtime is dominated by the inter-robot alignment module (a one-time computation for each robot), and the $3$DGS map optimizer (an on-going computation to integrate streaming data into the map). The runtime of the alignment module is dominated (90% - 99.5% of the runtime) by the localized SfM procedure. Each SfM alignment requires about 33.6s to solve. The remaining runtime is primarily the feature-matching verification for proposed image correspondences (approximately 2.6s). In comparison, computing the batch COLMAP solution on all images for the Oracle baseline takes between 2 and 5 hours depending on the configuration of COLMAP, illustrating the need for HAMMER’s alignment module, which requires a one-time computation (per device) of about 36 sec. For the continual $3$DGS map optimizer, processing a single data-frame ranges between 15 ms (66 Hz) when the map is sparsely populated to 38 ms (26 Hz). Slower iterations typically occur immediately after a robot has been aligned and all of its cached data is consolidated into the map.
Bandwidth and communication overhead required by HAMMER are dependent on sensor configuration. Data streams for each device in our hardware experiments operate at 30-35 Mbps. Of that total bandwidth, the depth data is 25-30 Mbps and the compressed color images are approximately 2 Mbps. This bandwidth can be sustained by high-end commercial WiFi routers. Applying depth compression [39] would substantially reduce the required bandwidth, although this technology was not directly supported by our devices. Down-sampling data in both transmission frequency and resolution could also reduce bandwidth overhead.
Methods Metrics Off-0 Apt-0 Apt-1 Apt-2 Avg.
CP-SLAM PSNR [dB] $\uparrow$ 28.56 26.12 12.16 23.98 22.71
[14] SSIM $\uparrow$ 0.87 0.79 0.31 0.81 0.69
LPIPS $\downarrow$ 0.29 0.41 0.97 0.39 0.51
Depth L1 [cm] $\downarrow$ 2.74 19.93 66.77 2.47 22.98
Time [min] $\downarrow$ 200+ 200+ 200+ 200+ 200+
MAGiC-SLAM PSNR [dB] $\uparrow$ 38.53 35.87 28.57 30.01 33.24
[13] SSIM $\uparrow$ 0.98 0.97 0.93 0.94 0.95
LPIPS $\downarrow$ 0.07 0.21 0.17 0.19 0.16
Depth L1 [cm] $\downarrow$ 0.48 0.72 5.85 1.17 2.05
Time [min] $\downarrow$ 85 146 158 144 133
HAMMER PSNR [dB] $\uparrow$ 43.02 40.89 31.43 33.78 37.28
SSIM $\uparrow$ 0.98 0.97 0.90 0.94 0.95
LPIPS $\downarrow$ 0.04 0.06 0.13 0.11 0.09
Depth L1 [cm] $\downarrow$ 0.94 1.27 2.40 1.25 1.47
Time [min] $\downarrow$ 8 8 8 8 8
TABLE I: Map reconstruction performance based on training view synthesis between HAMMER and other multi-robot mapping methods on ReplicaMultiAgent. Best and second best values are highlighted in green and yellow, respectively. CP-SLAM metrics are drawn from [13], with time estimated based on hardware used to run other methods.
IV-B4 Multi-robot Semantic Motion Planning
@@
As a brief qualitative case study and to highlight the utility of the semantic $3$DGS map for downstream tasks, we demonstrate that a map constructed online with HAMMER can be used for multi-robot motion planning task with language-specified goals in Basement after about 4 minutes of wall-clock time. Goals are determined by searching the map for the Gaussians with the highest semantic similarity to the query (e.g. “couch”). Once the goal is located, we use Splat-Nav [7] to plan a collision-free trajectory for each robot to the goal location. Figure 6 shows optimized trajectories for each of the four robots in Basement with 9 semantically-specified goals. Goal location required about 20ms, and the subsequent trajectories were computed within 0.5s. All robots successfully completed the task.
V Conclusion and Limitations
HAMMER has several limitations. First, devices without pose estimates cannot be used. Moreover, very noisy estimates or significant drift cannot be rectified by HAMMER’s pose refinement, resulting in degraded maps. Furthermore, the current implementation of HAMMER does not account for RGB-only devices, but we believe HAMMER can be extended to RGB-only by using monocular depth estimation. Finally, very short robot deployments render the one-time alignment cost prohibitively expensive.